Circular No. 7372

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/2000 D2 (LINEAR) An apparently asteroidal object discovered by LINEAR and posted on

the NEO Confirmation Page has been observed to be cometary by F. B. Zoltowski (Edgewood, NM; very diffuse image on Feb. 28.1 UT; 12′′ tail in p.a. 270^{o} on Mar. 1.1) and by C. Hergenrother (Catalina 1.54-m reflector; 8′′ coma and very faint 15′′ tail in p.a. 105^{o} on Mar. 1.3). Full astrometry and parabolic orbital elements (T = 2000 Mar. 2.3 TT, q = 2.323 AU, i = 156^{o}.9) appear on MPEC 2000-E04.

2000 UT  α_{2000}  δ_{2000}  m_{2}
Feb. 25.19852  6^{h}48^{m}19.^{s}35  +52^{o}53′53′′.6  18.1

SUPERNOVA 2000L IN UGC 5520 T. Puckett, Mountain Town, GA, reports the discovery of an apparent

supernova (mag 17.8) on an unfiltered CCD frame (limiting mag 19.3) taken with the Puckett Observatory 0.35-m supernova patrol telescope on Feb. 29.13 UT. The candidate is located at α = 10^{h}15^{m}01^{s}.82, δ = +65^{o}08′03′′.5 (equinox 2000.0), which is 6′′.9 east and 19′′.3 south of the center of UGC 5520. The new object was also present on an unfiltered CCD frame taken on Mar. 1.02, but it was not present on a frame taken on 2000 Jan. 7.25, and it does not appear on Palomar Sky Survey images taken on 1997 Jan. 11 (limiting mag ∼ 21.2) and 1953 Feb. 8 (limiting mag ∼ 20.1).

V4633 SAGITTARII Y. Lipkin and E. M. Leibowitz, Wise Observatory, Tel Aviv University,

report: “We performed CCD photometry of V4633 Sgr with the Wise Observatory 1-m telescope through an I filter during 36 nights in 1999. Timeseries analysis of the data reveals two distinct periodicities at 0.128791 ± 0.000004 and 0.125573 ± 0.000005 day. The full amplitudes of the periodicities are 0.097 and 0.081 mag, respectively. We cannot rule out that the 1-day aliases of these periodicities, at 0.147833 and 0.143674 day, are the true periods. The former alias is, to within the accuracy of our measurments, one of the periods reported in 1998 by Lipkin et al. (IAUC 6963). After re-examining the data of our six observing nights in 1998, we find that the second periodicity of 1999 is present in 1998 as well, with the same period as in 1999, to within the accuracy of our measurments. The full amplitudes of the periodicities in 1998 are 0.026 and 0.020 mag, respectively.”

2000 February 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green